OpenStudy (anonymous):

solve for the indicated variable 3x - 8y = 24 ; for y

OpenStudy (ajprincess):

1.Add 8y both sides 2.Subtract 24 both sides 3.Divide by 8 both sides

OpenStudy (anonymous):

3x-8y=24 -3x -3x __________________________ -8y= 24-3x -8y/-8 = 24-3x/-8 y= ??

OpenStudy (anonymous):

should my answer be .375x - 3 = y?

OpenStudy (ajprincess):

yup u r right. But it wud be really bettter if u give the answer as \[\frac{ 3 }{ 8 }x-3=y\]

OpenStudy (anonymous):

okay,thanx,ive another similiar one as well if you help Rx +sy = t ; for y (so =/ 0)

OpenStudy (anonymous):

(s =/ 0)

OpenStudy (ajprincess):

1.Subtract Rx both sides 2.Divide s both sides

OpenStudy (anonymous):

if a fixed cost is 55,000 v.c is 1.70 and s.p is 16 is ure answer 3.595?

OpenStudy (ajprincess):

vat is v.c and s.p?

OpenStudy (anonymous):

variable cost and selling price,i should of explained that i have to find out how many units to sell to break even

OpenStudy (ajprincess):

I am sorry. I hav no idea:(

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x/6 + 1/2 = 5/12?